. What is the temperature of a system that consists of 4.00 moles of argon behaving as an ideal gas with an internal energy of 15500 J?
1
Expert's answer
2020-10-14T14:05:06-0400

Solution.

Argon has three degrees of freedom, since its molecule consists of a single atom.

"\Delta U = \frac{3}{2} n \times R \times T"

"T = \frac{\Delta U}{\frac{3}{2} \times n \times R}"

T = 310.9 K

Answer:

T = 310.9 K